A boat travels 28 km upstream in 7 hours. It travels 48 km downstream in 4 hours. Find the speed of the boat and the speed of the stream.
Answer by ankor@dixie-net.com(22740)

A boat travels 28 km upstream in 7 hours.
It travels 48 km downstream in 4 hours.
Find the speed of the boat and the speed of the stream.
:
let s = speed of boat in still water
let c = speed of the current
then
(s-c) = speed up stream
(s+c) = speed down stream
:
Write an distance equation for each way. Dist = time * speed
7(s-c) = 28
4(s+c) = 48
Simplify these equations, divide the 1st by 7 and the 2nd by 4
s - c = 4
s + c = 12
---------------addition eliminates c find s
2s + 0 = 16
s = 16/2
s = 8 km/h boat speed in still water
then
8 - c = 4
-c = 4 - 8
-c = -4
c = 4 km/hr is the speed of the current
:
:
Check this in the original equations
7(8-4) = 28
4(8+4) = 48